eatsalad:

Solve the system of equations. 10x+3y+2z=5 7x+6y+2z=20 4x−3y+3z=7 Can someone pls show me step by step

SmokeyBrown:

Here is what I would like to try Combine the first and second equations with subtraction in order to eliminate the "z" terms: 10x+3y+2z=5 -(7x+6y+2z=20) ________________ 3x -3y = -15 Then, divide both sides of this resulting equation by 3 to get: x - y = -5 Add y to both sides to get the relationship: x = y - 5 Now we know that x is 5 less than y, and we've gotten one step closer to solving for x, which in turn will allow us to solve for y and z. Perhaps this will give you some idea of the methods you can use to go about this kind of problem. I won't solve the entire thing for you, but I'll be happy to give more guidance if needed. Please let me know if anything I demonstrated was confusing or if you need further help :)

eatsalad:

ok here is where i mess up u chose the first and second equation, why not the first and third?

eatsalad:

like, does it matter which i start first

eatsalad:

and why did u subtract i thought we always had to add

SmokeyBrown:

@eatsalad wrote:
like, does it matter which i start first
It does not necessarily matter which equations you choose to combine, since the result will always be the same. I chose to use the first and second to start because I thought they would give a neater result which would be easier to work with. Taking the first and third equations and combining them with addition would give: 10x+3y+2z=5 4x−3y+3z=7 ______________ 14x + 5z = 12 Actually, we can use this in combination with the results we found earlier (x = y - 5) By replacing the x in this equation with (y-5) we get 14(y-5) + 5z = 12 Which simplifies to 14y - 70 + 5z = 12 Then 14y + 5z = 82 ... Which gives a relationship between y and z. Although this is not a very neat relationship, it could still be another hint toward solving for x, y, and z... But maybe it would be more helpful to go in another direction? Honestly, I'm not sure if there is a systematic way to solve these kinds of problems without using matrices (that may be a bit too advanced for now), so my best advice is to keep trying things until you happen upon something that works ':D

eatsalad:

thank you i'll keep trying , im in 11th grade and i think this year we'll start learning matrices, but this was just a review to begin

SmokeyBrown:

No problem. I'm willing to bet there's other users around who are more familiar with this topic than I am. Maybe @mhanifa or @florisalreadytaken ? They tend to be rather active in the Mathematics section. In any case, I'm sure you'll be able to get the help you need :)

SmokeyBrown:

I think I have another idea. Take the relationship (x = y - 5), which we got by combining the first and second equations. Use that to replace x with (y-5) in the third equation: 4x−3y+3z=7 Becomes 4(y-5) -3y + 3z = 7 Distribute the 4 over y and -5 to get 4y - 20 - 3y + 3z = 7 Combine the common y terms: -20 + y + 3z = 7 Add 20 to both sides: y + 3z = 27 Subtract 3z from both sides: y = 27 - 3z Now we have a relationship between y and z. Use this together with the relationship we found earlier between x and y, and we now have a way to substitute terms so that any equation will have only 1 variable. In this way, we should be able to solve for "z". After that, we can solve for x and y.

SmokeyBrown:

I shall demonstrate using the first equation (though any of them would work: 10x+3y+2z=5 Use (x = y-5) to substitute x with (y-5): 10(y-5) + 3y + 2z = 5 Distribute the 10 to y and -5: 10y - 50 + 3y + 2z = 5 Add the common y terms: 13y - 50 + 2z = 5 Add 50 to both sizes: 13y + 2z = 55 Use (y = 27 - z) to substitute y for (27 - z) 13(27 - z) + 2z = 55 Distribute the 13 to 27 and z: 351 -13z + 2z = 55 Combine the common z terms: 351 - 11z = 55 Subtract 351 from both sides: -11z = -296 Divide both sides by -11: z = 296/11 And just like that, we have the value of z! From here, finding the values of x and y should be a piece of cake, especially given those relationships we found earlier, (x = y-5) and (y = 27-z) I trust I can leave that to you?

eatsalad:

I think i did something really wrong i got mine a different way (the way my lesson said) where i began with the first and last equation and for z I got 8

SmokeyBrown:

@eatsalad wrote:
I think i did something really wrong i got mine a different way (the way my lesson said) where i began with the first and last equation and for z I got 8
Ah, then perhaps I made an error in my calculations. That wouldn't be too surprising =D Would you mind going through the steps you took to get that result, so we can check?

eatsalad:

i did from the begginng 10x+3y+2z=5 -- 4x+-3y+3z=7 i aded the first and third to eliminate y and got 14x+5z=12 Then 7x+6y+2z=20 and added 8x+-6y+6z=14 i got that by Multiplying it by 2 so that the y term then i got 15x+8z=34 14x+5z=12 15x+8z=34 112x+40z=96 −75x−40z=−170 get that bymultiplying first equation by 8 and the second equation by −5 then here is where i get x, y and z 112x+40z=96 −75x−40z=−170 i added the equation and got 37x=74=x=-2 now 15(−2)+8z=34 =z=8 10(−2)+3y+2⋅8=5 and got y=3

eatsalad:

but thats a really long tiring way

SmokeyBrown:

Hey, that's awesome! I didn't see any errors in your work, so I'm pretty confident you got the correct answer. Well done! It does make me a little curious about what I did wrong, but that's not as important. I'm glad you were able to figure it out for yourself

@eatsalad wrote:
but thats a really long tiring way
On the contrary, I thought your explanation made a lot of sense. Good work :)

SmokeyBrown:

I think I caught my error. Back in this step, I substituted y with (27-z), but it should have been (27 - 3z):

@smokeybrown wrote:
I shall demonstrate using the first equation (though any of them would work: 10x+3y+2z=5 Use (x = y-5) to substitute x with (y-5): 10(y-5) + 3y + 2z = 5 Distribute the 10 to y and -5: 10y - 50 + 3y + 2z = 5 Add the common y terms: 13y - 50 + 2z = 5 Add 50 to both sizes: 13y + 2z = 55 Use (y = 27 - z) to substitute y for (27 - z) 13(27 - z) + 2z = 55 Distribute the 13 to 27 and z: 351 -13z + 2z = 55 Combine the common z terms: 351 - 11z = 55 Subtract 351 from both sides: -11z = -296 Divide both sides by -11: z = 296/11 And just like that, we have the value of z! From here, finding the values of x and y should be a piece of cake, especially given those relationships we found earlier, (x = y-5) and (y = 27-z) I trust I can leave that to you?
If I hadn't made this typo, the calculation would have been z = 296/37, which equals 8, as you found. Personally, I like your method better. But, since we both (eventually) got the right answer, I feel even more confident that your results are correct! Once again, well done!

eatsalad:

Thank you for your time

SmokeyBrown:

@eatsalad wrote:
Thank you for your time
And you for your patience :)